Rachel spent $16.18 for some cans of dog food cost in 79 cents each and some cans of cat food costing 69 cents each. She bought two more cans of cat food than of dog food. How many cans of each did she buy?
Answer by stanbon(75887)

Rachel spent $16.18 for some cans of dog food cost in 79 cents each and some cans of cat food costing 69 cents each. She bought two more cans of cat food than of dog food. How many cans of each did she buy?
--------------------------
Let the number of cans of dog food = "x".
---
Then the number of cans of cat food = "x+2".
---
Value equation:
value + value = value
79x + 69(x+2) = 1618 cents
---
79x + 69x + 138 = 1618
148x = 1480
x = 10 (# of cans of dog food)
----
x+2 = 12 (# of cans of cat food)
==================================
Cheers,
Stan H.